The motion of the particle executing shm is given by x=0.01sin100pi(t+.05) where x in metres and time is in seconds The time period is

Grade:12

1 Answers

Rohith Gandhi
14 years ago

Dear Pallavi,

x = 0.01sin(100pi*t + 5*pi) => x =  -0.01sin(100*pi*t)

 

hence from equation, omega = angular frequency = 100*pi = 2*pi/T

 

=>Time period ,  T = 2/100 = 0.02 seconds


Please feel free to post as many doubts on our discussion forum as you can. If you find any question
Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We
are all IITians and here to help you in your IIT JEE preparation.


All the best Pallavi !!!




Regards,

Askiitians Experts
Rohith Gandhi P